OpenStudy (anonymous):

solve. -2y^(2)-16y≥24

OpenStudy (athe):

\[-2y^2-16y \ge 24 \rightarrow y^2+8y+12 \le 0 \rightarrow (y+2)(y+6) \le 0 \rightarrow y \in \left[ -6, -2 \right]\]